# Pilot-Quantum: A Quantum-HPC Middleware for Resource, Workload and Task Management

### v2

@@ III-B System Architecture
The Pilot Abstraction’s multi-level scheduling suits quantum-HPC environments for two reasons: (i) heterogeneity and complexity of quantum and classical applications complicate static workload and resource assignment, and (ii) early quantum era co-design of applications and hardware requires integrated application-level resource management. While small applications can use hard-wired integration, larger hybrid systems need more adaptive resource management. Hence, hierarchical resource management capabilities are crucial for developing and operating quantum-enabled workflow systems efficiently. Next, we present each layer of the system.
@@ III-B1 L4 Workflow Layer and Abstractions:
For (ii), applications must decompose their problems into a set of independently executable tasks, collectively forming the workload. Tasks can encapsulate units of quantum computations (e. g., circuit executions) and classical computations (e. g., pre- and post-processing). While early quantum applications require high control over problem decomposition and workload, Pilot-Quantum is expected to interface with higher-level workflow systems and application libraries.
Tasks are defined using a task_description and submitted via the submit_task API. Pilot-Quantum also integrates with native abstractions of different frameworks, e. g., Dask and Ray. Tasks can be assigned: (i) directly to a specific pilot or (ii) to a group of pilots. For (ii), the pilot manager assigns tasks to pilots, optimizing throughput by load balancing across all pilots. Please, refer to [78] for API details.
@@ III-B2 L3/L2 Workload and Task Layers:
After receiving the pilot_description, the Pilot-Manager allocates requested resources via the local resource management system, (steps 1-3 in Figure 2). In HPC environments, a placeholder job, which starts the Pilot-Agent is queued via Slurm. Once active, the Pilot Agent manages the allocated resources and executes tasks either directly or through external execution engines. External execution engines are supported via plugins; currently, Pilot-Quantum supports Dask [79] and Ray [68]. Further, it integrates with low-level runtimes (e. g., CUDA and MPI). Quantum hardware access is handled via classical tasks that utilize, e. g., Qiskit Providers or PennyLane plugins.
The choice of an external execution engine depends on workload characteristics and resource infrastructure. Dask is well-suited for data-intensive workloads, such as data-parallel streaming and machine learning tasks [79]. Pilot-Quantum extends its capabilities with a plugin for Ray [68], a versatile distributed computing framework that facilitates seamless integration of GPUs (essential for quantum simulations) and custom resources like QPUs. For tightly coupled HPC tasks, such as distributed state vector simulations, the MPI and CUDA launch mechanisms are crucial. The agent leverages Slurm’s srun command to efficiently distribute tasks across nodes, cores, and GPUs, ensuring optimal parallel execution considering machine architecture and network topology.
@@ III-B3 L1 Resource Layer:
As described, Pilot-Quantum aims to abstract heterogeneous resources through its plugin mechanism for various runtimes. In addition, it can be integrated with various resource-specific tools, e. g., QPU-specific transpilers.
@@ IV Evaluation and Results
This section characterizes Pilot-Quantum’s capabilities using the Quantum Mini App framework [19, 76]. Mini-apps are simplified versions of quantum-HPC applications, representing key characteristics and performance bottlenecks through recurring execution patterns (motifs). We extend the circuit simulation mini-app [76] and develop a QML mini-app.
@@ IV-A Pilot-Quantum Characterization
TABLE I: Average Throughput and Duration by Task Size
@@ IV-B Circuit Simulation Mini-App
We extend the circuit simulation mini-app [19] with quantum hardware support (section IV-B1), an implementation of distributed state vector simulation (section IV-B2), and a circuit cutting motif (section IV-B3).
@@ IV-B1 Circuit Execution:
Figure 3: Circuit Execution on IonQ, IBM Eagle and Qiskit Aer (CPU and GPU): Comparing the execution times for 1,024 random quantum circuits (2 to 28 qubits) on IonQ Quantum Cloud and Qiskit’s Aer simulators (CPU and GPU) using Pilot-Quantum (Ray) on one Perlmutter node. For IBM Eagle, we only executed eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the 1,024-circuit experiments. IonQ required 4,260 seconds for 28 qubits, while Aer (CPU) took 469.7 seconds, and Aer (GPU) completed it in just 49 seconds. The scaled execution time of the IBM Eagle QPU is 199 seconds.
We evaluate various hardware and simulated QPU backends, specifically: (i) in-process simulators running on application-allocated resources, like Qiskit Aer [81], (ii) cloud simulators, and (iii) a QPU via the IBM cloud. For (i), we investigate Qiskit’s Aer CPU and GPU simulators. For (ii), we investigate IonQ’s cloud simulator [82]. Further, we investigate IBM’s Eagle QPU hardware accessed via the cloud. We used Pilot-Quantum (Ray) on one Perlmutter node to manage the execution of $1024$ random quantum circuits with $2$-$28$ qubits across these backends.
@@
For IBM Eagle, we executed only eight random circuits for each qubit configuration and scaled their mean execution time to be comparable with the $1,024$-circuit experiments. In contrast to the simulation experiments, the execution time varies slightly but does not increase exponentially.
IV-B2 Distributed State Vector Simulation:
enables efficient simulation of large quantum circuits with higher qubit counts and depth by using multiple nodes and GPUs [83, 84]. The distributed state vector motif [19] uses PennyLane’s lightning.gpu [85] v0.35.0. This device is a quantum circuit simulator implemented on NVIDIA’s cuQuantum [46], using a C++ backend and CUDA/MPI for distributed computations and data movements.
@@ IV-B2 Distributed State Vector Simulation:
Figure 4: PennyLane lightning.gpu Distributed State Vector Simulations: Computing the expectation value of a 2-layer strongly entangling layered (SEL) circuit with and without gradient calculation using PennyLane’s lightning.gpu device on 256 A100 80 GB GPUs on Perlmutter.
Figure 4 shows distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86]. We compare circuit execution with and without gradient computation, using adjoint differentiation [87] for gradients, which offers lower memory and computational overhead than alternatives like finite difference requiring multiple circuit evaluations. We investigated $31$-$37$ qubits on 256 GPUs, the upper qubit range limited by gradient computation overheads, without which can scale to $41$ qubits. Results indicate GPU resources must be balanced between simulation and other intensive tasks like ML for optimal performance.
IV-B3 Circuit Cutting
[88] enables running larger circuits on quantum hardware at the cost of increased sampling and classical processing overhead. It decomposes larger circuits into smaller ones, referred to as subcircuits, by cutting wires and gates, allowing execution within hardware and simulator constraints. However, the number of required subcircuit executions increases exponentially with circuit size and number of cuts [89]. The sampling results from the subcircuits are combined to reconstruct the original circuit outcome.
This circuit-cutting motif uses Qiskit’s circuit cutting addon [90]. It uses Qiskit’s EfficientSU2 ansatz, commonly used in many variational algorithms, comprising single qubit rotations and linear CNOT gates between adjacent qubits. Qiskit’s expectation value reconstruction combines results from smaller sub-experiments. It provides options to run full circuit and/or circuit cutting simulations, enabling comparison of the error induced by circuit cutting. It gathers comprehensive performance metrics throughout the process.
@@ IV-B3 Circuit Cutting
In the following experiments, we use a $34$-qubit quantum circuit of the described ansatz. For circuit cutting, we utilize two configurations: (i) subcircuit size $17$ and (ii) subcircuit size $12$. We used the find_cuts method to determine the best cut and generate necessary subcircuit executions: $72$ for scenario (i), and approximately 2,100 subcircuit experiments for (ii). We compare these circuit-cutting scenarios to a full circuit simulation using Qiskit’s Aer distributed state vector simulator (scenario (iii)) with multi-node GPU and MPI support [81, 91]. All experiments are repeated at least three times.
Figure 5 illustrates performance scaling from $1$ to $16$ nodes ($4$ to $64$ GPUs). In scenario (i), with fewer subexperiments, optimal performance is achieved on a single node. Conversely, scenario (ii) shows the best performance with $2$ nodes, achieving a $\sim 1.5$ speedup, demonstrating that additional resources executed the 2,100 subexperiments faster. When comparing to the sequential runtimes on 1 core for (i) 58 sec and (ii) 304 sec, we were able to achieve a (i) 2.1×2.1$\times$2.1 × and (ii) 3.5×3.5$\times$3.5 × speedup. Scenario (iii), the distributed state vector simulation, outperforms both circuit-cutting scenarios when using more than $4$ nodes. The best performance was achieved with 16 nodes, showing a $\times5$ speedup.
In summary, different circuit-cutting strategies lead to varying optimal performance points, highlighting the importance of application-level scheduling, as enabled by Pilot-Quantum, for efficient resource and task management. Hence, an adaptive execution strategy combining multiple simulation methods (e. g., circuit cutting and distributed state vector simulation) and leveraging heterogeneous resources (e. g., CPUs, GPUs, and QPUs) will be crucial for effective application scaling.
@@ V-A Discussion
This section qualitatively compares Pilot-Quantum with Qiskit Serverless, Covalent, and CUDA-Q (see section  II-D) by mapping each across the four architectural layers (see Figure 8 and Table II).
L4: Qiskit Serverless and Pilot-Quantum use task-based models, while Covalent provides workflow-level abstraction at $L4$ layer. CUDA-Q’s kernel-based approach supports tighter integration modes (e. g., error correction) and large-scale computational demands (e. g., distributed state vector simulations), and can be used in conjunction with Pilot-Quantum.
While each framework supports workflow dynamism, i. e., enabling new task creation based on prior results, including quantum tasks, only Covalent and Pilot-Quantum allow adding resources at runtime. At the application-level, Qiskit Serverless is designed to integrate with Qiskit’s libraries, while Covalent and Pilot-Quantum are framework-agnostic. In contrast, CUDA-Q is optimized for tight coupling of quantum and classical tasks.
L3: Both Covalent and Pilot-Quantum offer high-level control over resource allocation and task management. Qiskit Serverless allows application-level task management but not resource allocation, while CUDA-Q, uses a fixed HPC resource model with MPI and CUDA for task execution, lacking resource management capabilities and requiring manual integration with resource management systems like Slurm.
L2: At the task execution layer, most frameworks use task-based systems like Dask and Ray for parallel execution across diverse resources, decoupling application from execution details (e. g., scheduling). They allow some application-level control of task and resource mapping for optimally utilizing accelerator and data parallelism. Resource mapping depends heavily on application details, such as the amount of data per task, making application-level scheduling crucial for optimal performance. CUDA-Q, with its runtime, directly manages kernel offloading to QPUs or accelerators. While supporting some dynamism, CUDA-Q’s compiled nature makes it less flexible than other task execution systems.
@@
L1: At the $L1$ resource layer, Qiskit Serverless, designed as a managed serverless cloud service, offers limited high-level control, e.g.., the precise number of cores, memory, and GPUs cannot be specified. It integrates well with resource-specific compilers and runtimes. Covalent also provides a managed serverless service but lacks quantum-specific integrations. Both frameworks support deployment in dedicated compute environments, typically requiring manual setup with tools like Terraform. CUDA-Q does not include a cloud or cluster resource manager and assumes ownership of resources.
Summary and Limitations: Each of the above systems targets different scenarios and quantum-HPC integration modes: Qiskit Serverless is workload-centric, Covalent focuses on workflows, CUDA-Q aims for tighter integration (HPC-for-Quantum), while the others prioritize scaling loosely coupled workloads (Quantum-x-HPC modes). Resource, workload, and task management are secondary for most systems. Pilot-Quantum addresses this gap by enabling seamless integration with heterogeneous resources, simplifying deployment, and supporting flexible resource use at the application level, crucial for diverse workloads [19].
@@ V-B Conclusion and Future Work
Pilot-Quantum is middleware that enhances hybrid quantum-classical workload management on heterogeneous infrastructure. Based on analysis of execution patterns and motifs [19], it optimizes resource allocation and task execution for efficient use of heterogeneous resources. Pilot-Quantum is designed to handle complex hybrid applications with loosely coupled parallelism at our conceptual architecture’s workload layer ($L3$). In addition, it can integrate with tightly coupled application components, e. g., MPI/CUDA-based distributed state vector simulations. Our experimental results demonstrate Pilot-Quantum capabilities with significant improvements in task execution times, e. g., a 3.5×3.5$\times$3.5 × speedup for circuit cutting and a 15×15$\times$15 × speedup for QML. Pilot-Quantum will be a critical tool for quantum-HPC workload and task management, addressing resource and application heterogeneity as quantum computing evolves.
In the future, we plan to support Pilot-Quantum with higher-level frameworks like Covalent and enable direct quantum resource access via OpenQASM and QIR-based tasks. We will also investigate scheduling algorithms for heterogeneous resource task assignments. A key step will be developing Q-Dreamer, a predictive tool for estimating optimal resources for specific workloads and dynamically scheduling tasks based on real-time availability and system performance.

### v3

@@ III-B System Architecture
The Pilot Abstraction’s multi-level scheduling suits quantum-HPC environments for two reasons: (i) heterogeneity and complexity of quantum and classical applications complicate static workload and resource assignment, and (ii) early quantum era co-design of applications and hardware requires integrated application-level resource management. While small applications can use hard-wired integration, larger hybrid systems need more adaptive resource management. Hence, hierarchical resource management capabilities are crucial for efficiently developing and operating quantum-enabled workflow systems. Next, we present each layer of the system.
@@ III-B1 L4 Workflow Layer and Abstractions:
For (ii), applications must decompose their problems into independently executable tasks, collectively forming the workload. Tasks can encapsulate units of quantum computations (e. g., circuit executions) and classical computations (e. g., pre- and post-processing). While early quantum applications require high control over problem decomposition and workload, Pilot-Quantum is expected to interface with higher-level workflow systems and application libraries.
Tasks are defined using a task_description and submitted via the submit_task API. Pilot-Quantum also integrates with native abstractions of different frameworks, e. g., Dask and Ray. Tasks can be assigned (i) directly to a specific pilot or (ii) to a group of pilots. For (ii), the pilot manager assigns tasks to pilots, optimizing throughput by load balancing across all pilots. Please refer to [78] for API details.
@@ III-B2 L3/L2 Workload and Task Layers:
After receiving the pilot_description, the Pilot-Manager allocates requested resources via the local resource management system (steps 1-3 in Figure 2). In HPC environments, a placeholder job, which starts the Pilot-Agent is queued via Slurm. Once active, the Pilot Agent manages the allocated resources and executes tasks directly or through external execution engines. External execution engines are supported via plugins; currently, Pilot-Quantum supports Dask [79] and Ray [68]. Further, it integrates with low-level runtimes (e. g., CUDA and MPI). Quantum hardware access is handled via classical tasks that utilize, e. g., Qiskit Providers or PennyLane plugins.
The choice of an external execution engine depends on workload characteristics and resource infrastructure. Dask is well-suited for data-intensive workloads, such as data-parallel streaming and machine learning tasks [79]. Pilot-Quantum extends its capabilities with a plugin for Ray [68], a versatile distributed computing framework that facilitates seamless integration of GPUs (essential for quantum simulations) and custom resources like QPUs. The MPI and CUDA launch mechanisms, such as distributed state vector simulations, are crucial for tightly coupled HPC tasks. The agent leverages Slurm’s srun command to efficiently distribute tasks across nodes, cores, and GPUs, ensuring optimal parallel execution considering machine architecture and network topology.
@@ III-B3 L1 Resource Layer:
As described, Pilot-Quantum aims to abstract heterogeneous resources through its plugin mechanism for various runtimes. In addition, it can be integrated with resource-specific tools, e. g., QPU-specific transpilers.
@@ IV Evaluation and Results
This section characterizes Pilot-Quantum’s capabilities using the Quantum Mini-App framework [22, 76]. Mini-apps are simplified versions of quantum-HPC applications, representing key characteristics and performance bottlenecks through recurring execution patterns (motifs). We extend the circuit simulation mini-app [76] and develop a QML mini-app.
@@ IV-A Pilot-Quantum Characterization
TABLE I: Average Throughput and Duration by Task Size
@@ IV-B Circuit Simulation Mini-App
We extend the quantum simulation mini-app [22] by adding and evaluating different cloud backends (incl. physical quantum hardware) (section IV-B1), an implementation of distributed state vector simulation (section IV-B2) and circuit cutting (section IV-B3).
@@ IV-B1 Circuit Execution:
Figure 3: Circuit Execution on IonQ, IBM Eagle and Qiskit Aer (CPU and GPU): Comparing the execution times for 1,024 random quantum circuits (2 to 28 qubits) on IonQ Quantum Cloud and Qiskit’s Aer simulators (CPU and GPU) using Pilot-Quantum (Ray) on one Perlmutter node. We only executed eight random circuits for each qubit configuration for IBM Eagle and scaled their mean execution time to be comparable with the 1,024-circuit experiments. IonQ required 4,260 seconds for 28 qubits, while Aer (CPU) took 470 seconds, and Aer (GPU) completed it in just 49 seconds. The scaled execution time of the IBM Eagle QPU is 199 seconds.
We evaluated various quantum backends, including (i) in-process simulators using Qiskit Aer [81], (ii) IonQ cloud simulators accessed via Qiskit IonQ provider [82], and (iii) IBM’s Eagle QPU via IBM cloud using Qiskit’s Runtime library. We deployed Pilot-Quantum (Ray) on a Perlmutter node to manage the execution of 1,024 random circuits across these backends, ranging from $2$ to $28$ qubits.
@@
To demonstrate Pilot-Quantum’s capabilities in executing quantum circuits in parallel on physical quantum hardware, we assume that 256 QPUs are available with exclusive access (i.e., no queuing time). Due to quantum hardware costs and limited availability, we execute only eight random circuits instead of 1024. The average execution time for these circuits is then scaled by a factor of 4 ($1,024/256=4$) to estimate the performance for 256 QPUs processing 1,024 random circuits. Additionally, the time accounts for the pilot quantum compute overhead of 2 seconds for 1024 circuits on a single node with 256 tasks. In contrast to the simulation experiments, the execution time varies slightly but does not increase exponentially.
IV-B2 Distributed State Vector Simulations
enables efficient simulation of large quantum circuits with higher qubit counts and depth using multiple nodes and GPUs [83, 84]. It is based on PennyLane’s lightning.gpu [85] (v0.41.0), which leverage a cuQuantum [24] for distributed simulation. We use $64$ nodes and $256$ GPUs.
@@ IV-B2 Distributed State Vector Simulations
Figure 4: PennyLane lightning.gpu Distributed State Vector Simulations: Computing the expectation value of a 2-layer strongly entangling layered (SEL) circuit with and without gradient calculation using PennyLane’s lightning.gpu device on 256 A100 40 GB GPUs on Perlmutter.
Figure 4 shows distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86] using up to $39$ qubits. We compare circuit execution (i) with and (ii) without gradient calculation. For (i), we use adjoint differentiation [87], which reduces memory and compute overhead compared to finite difference methods. For (i), we could run circuits of up to $34$ qubits, and for (ii), for up to $39$ qubits. The results underscore resource management’s importance, including careful memory partitioning for state vectors and gradient data.
Figure 4 shows the distributed state vector simulation of a 2-layer strongly entangling layered (SEL) circuit [86], which is frequently utilized for classification tasks. We compare circuit execution with and without gradient computation. We employ adjoint differentiation [87] for gradient calculation, a method optimized for quantum simulations that offers lower memory and computational overhead compared to alternatives like finite difference, which necessitates multiple circuit evaluations. We investigate qubit counts from $30$ to $34$ using $256$ GPUs. The overhead of gradient calculation limits us to $34$ qubits on $256$ GPUs, while without gradient computation, we scale up to $39$ qubits. The results indicate that classical resources must be allocated appropriately between simulation and other computationally intensive tasks like ML to ensure optimal performance.
IV-B3 Circuit Cutting [88]
enables running larger circuits on quantum hardware at the cost of increased sampling and classical processing overhead. It decomposes larger circuits into smaller ones, referred to as subcircuits, by cutting wires and gates, allowing execution within hardware and simulator constraints. The sampling results from the subcircuits are combined to reconstruct the original circuit outcome. However, the number of required subcircuit executions increases exponentially with circuit size and number cuts [89].
This circuit-cutting motif uses Qiskit’s circuit-cutting addon [90]. It utilizes Qiskit’s EfficientSU2 ansatz, commonly used in many variational algorithms, comprising single qubit rotations and linear CNOT gates between adjacent qubits. Qiskit’s expectation value reconstruction combines results from smaller sub-experiments. It provides options to run full circuit or circuit cutting simulations, enabling comparison of the circuit cutting error. It gathers comprehensive performance metrics throughout the process.
@@ IV-B3 Circuit Cutting [88]
We use a $34$-qubit quantum circuit of the described ansatz in the following experiments. For circuit cutting, we utilize two configurations: (i) subcircuit size $17$ and (ii) subcircuit size $12$. We used the find_cuts method to determine the best cut and generate necessary subcircuit executions: $72$ for scenario (i), and approximately 2,100 subcircuit experiments for (ii). We compare these circuit-cutting scenarios to a full circuit simulation using Qiskit’s Aer distributed state vector simulator (scenario (iii)) with multi-node GPU and MPI support [81, 91]. All experiments are repeated at least three times.
Figure 5 illustrates performance scaling from $1$ to $16$ nodes ($4$ to $64$ GPUs). In scenario (i), with fewer subexperiments, optimal performance is achieved on a single node. Conversely, scenario (ii) shows the best performance with $2$ nodes, achieving a 1.5$\sim$ 1.5$\times$ 1.5 × speedup, demonstrating that additional resources executed the 2,100 subexperiments faster. When comparing to the sequential runtimes on one core for (i) 58 sec and (ii) 304 sec, we were able to achieve a (i) 2.1×2.1$\times$2.1 × and (ii) 3.5×3.5$\times$3.5 × speedup. The distributed state vector simulation (scenario (iii)) outperforms both circuit-cutting scenarios when using more than $4$ nodes. The best performance was achieved with 16 nodes, showing a $\times5$ speedup.
In summary, different circuit-cutting strategies lead to varying optimal performance points, highlighting the importance of application-level scheduling for efficient resource and task management, as enabled by Pilot-Quantum. Hence, an adaptive execution strategy combining multiple simulation methods (e. g., circuit cutting and distributed state vector simulation) and leveraging heterogeneous resources (e. g., CPUs, GPUs, and QPUs) will be crucial for effective application scaling.
@@ V-A Discussion
This section qualitatively compares Pilot-Quantum with Qiskit Serverless, Covalent, and CUDA-Q by mapping each across the four architectural layers (see Figure 8 and Table II).
L4: Qiskit Serverless and Pilot-Quantum use task-based models, while Covalent provides workflow-level abstraction at $L4$ layer. CUDA-Q’s kernel-based approach supports tighter integration modes (e. g., error correction) and large-scale computational demands (e. g., distributed state vector simulations). It can be used in conjunction with Pilot-Quantum.
While each framework supports workflow dynamism, i.e., enabling new task creation based on prior results, including quantum tasks, only Covalent and Pilot-Quantum allow adding resources at runtime. At the application level, Qiskit Serverless is designed to integrate with Qiskit’s libraries, while Covalent and Pilot-Quantum are framework-agnostic. In contrast, CUDA-Q is optimized for tight coupling of quantum and classical tasks.
L3: Both Covalent and Pilot-Quantum offer high-level control over resource allocation and task management. In contrast, Qiskit Serverless allows application-level task management but not resource allocation. At the same time, CUDA-Q uses a fixed HPC resource model with MPI and CUDA for task execution, but it lacks resource management capabilities and requires manual integration with resource management systems like Slurm.
L2: At the task execution layer, most frameworks use task-based systems like Dask and Ray for parallel execution across diverse resources, decoupling application from execution details (e.g., scheduling). They allow some application-level control of task and resource mapping for optimally utilizing accelerator and data parallelism. Resource mapping depends heavily on application details, such as the amount of data per task, making application-level scheduling crucial for optimal performance. The CUDA-Q runtime directly manages task execution on QPUs or accelerators. While supporting some dynamism, CUDA-Q’s compiled nature makes it less flexible than other task execution systems.
@@
L1: At the $L1$ resource layer, Qiskit Serverless, designed as a managed serverless cloud service, offers limited high-level control, e.g., the precise number of cores, memory, and GPUs cannot be specified. It integrates well with resource-specific compilers and runtimes. Covalent also provides a managed serverless service but lacks quantum-specific integrations. Both frameworks support deployment in dedicated compute environments, typically requiring manual setup with tools like Terraform. CUDA-Q does not include a cloud or cluster resource manager and assumes ownership of resources.
Summary and Limitations: Each of the above systems targets different scenarios and quantum-HPC integration modes: Qiskit Serverless is workload-centric, Covalent focuses on workflows, CUDA-Q aims for tighter integration (HPC-for-Quantum), while the others prioritize scaling loosely coupled workloads (Quantum-x-HPC modes). Resource, workload, and task management are secondary for most systems. Pilot-Quantum addresses this gap by enabling seamless integration with heterogeneous resources, simplifying deployment, and supporting flexible resource use at the application level, crucial for managing diverse workloads [22].
@@ V-B Conclusion and Future Work
Pilot-Quantum is middleware that enhances resource, workload, and task management in quantum-HPC systems. Pilot-Quantum is designed to handle complex hybrid applications with loosely coupled parallelism at our conceptual architecture’s workload layer ($L3$). In addition, it can integrate with tightly coupled application components, e. g., MPI/CUDA-based distributed state vector simulations. Our experimental results demonstrate Pilot-Quantum capabilities with significant improvements in task execution times, e.g., a 3.5×3.5$\times$3.5 × speedup for circuit cutting and a 15×15$\times$15 × speedup for QML. Pilot-Quantum will be a critical tool for quantum-HPC workload and task management, addressing resource and application heterogeneity as quantum computing evolves.
In the future, we plan to support Pilot-Quantum with higher-level frameworks like Covalent and enable direct quantum resource access via OpenQASM and QIR-based tasks. We also aim to investigate scheduling algorithms for heterogeneous resource task assignments. As a future work, a key step will be developing Q-Dreamer, a predictive tool for estimating optimal resources for specific workloads and dynamically scheduling tasks based on real-time availability and system performance.